OpenStudy (anonymous):

Medals will be given!! Please help!! If (x, 4) and (2, -1) are 2 points on a line and the slope is ½ , what is the value of x?

OpenStudy (amistre64):

the is a formula that is given that defines how to find the slope of a line between 2 points

OpenStudy (amistre64):

given 2 points (x,y) and (a,b) the slope between them can be calculated by:\[slope:\frac{y-b}{x-a}\]

OpenStudy (amistre64):

since we know all the values except for x, this fills in as: \[\frac12=\frac{4-(-1)}{x-2}\]

OpenStudy (anonymous):

then what?

OpenStudy (amistre64):

then you solve for x of course

OpenStudy (amistre64):

\[\frac12=\frac{4-(-1)}{x-2}\] \[\frac12=\frac{4+1}{x-2}\] \[\frac12=\frac{5}{x-2}\] cross multiply \[x-2=5(2)\]

OpenStudy (anonymous):

so x=12?

OpenStudy (amistre64):

x=12 is fine yes